1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
The Werewolf  Der Werwolf  Varulven
tradukita de Alexander Gross  tradukita de Sven Collberg

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“ En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der seine Pfoten geduldig kreuzte vor dem Toten: mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

�Der Werwolf“ – sprach der gute Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s nennt, den Werwolf, - damit hat’s ein End!“ “Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Dem Werwolf schmeichelten die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“ Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

Der Dorfschullehrer aber musste

gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

…  …  …

2/2

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

http://www.poezio.net/version?poem-id=825&version-id=,1657,1652,1654